OpenStudy (anonymous):

A rifle that shoots bullets at 460 m/s is to be aimed at a target 45.7 m away. If the center of the target is level with the rifle, how high above the target must the rifle barrel be pointed so that the bullet hits dead center?

OpenStudy (anonymous):

never mind i figured it out

OpenStudy (anonymous):

p.e=k.e mgh=1/2mv2 h=v2/2g h=(360)2/2x10 try??